OpenStudy (cicilybailey14):

I medal and Hug! The first five terms of a sequence are given below. 16 , 25 , 34 , 43 , 52 , ... Determine which of the following formulas gives the nth term of this sequence. A. 8 + 8n B. 25 - 9n C. 7 + 9n D. 24 - 8n

OpenStudy (cicilybailey14):

@ganeshie8 @Secret-Ninja

OpenStudy (cicilybailey14):

Okay :)

OpenStudy (cicilybailey14):

Here are my choice's :D A. n = (26 × 9) - 2 B. n = 26 × 29 C. n = 26 ÷ 9 D. n = 26 × 9

OpenStudy (anwaarullah):

its C,,,,,,,,,,,16+9=25 25+9=34 34+9= 43 and so on